# Scene-aware Human Motion Forecasting via Mutual Distance Prediction

### v3

@@ 4.2 Metrics, Baselines & Implementation
Baselines. We compare our method with three scene-aware human motion prediction models: ContAware [20], GIMO [40] and STAG [27]. ContAware [20] and STAG [27] are state-of-the-art methods which first predict the per-joint contact map and then the future motion conditioning on the predicted contact map. We train their model using the official implementation on our datasets, following their original settings to predict only human joints instead of SMPL-X poses. GIMO [40] is another state-of-the-art scene-aware human motion prediction method which mainly focuses on predicting long-term future poses given the past gaze information, and the frame rate of their motion is 2Hz. The original model of GIMO predicts future 10 frames (5 seconds) given the past 6 frames (3 seconds). We slightly adapt their official code to our settings for evaluation.
Implementation details. Our models are implemented in Pytorch [24] and trained using the ADAM [13] optimizer. Our mutual distance prediction model and motion forecasting model are first trained for 40 epochs separately and then fine-tuned together for 1 epoch. The original learning rate is set to 0.0005 and decays by half at the 30-th epoch for both models. The training on GTA-IM takes around 40 hours for each step on a 24GB NVIDIA RTX4090, respectively, while training on PROX dataset takes around 17 hours. For all datasets, the loss weights of the motion forecasting model ($\lambda_{1}$,$\lambda_{2}$,$\lambda_{3}$,$\lambda_{4}$) are set to $1.0,0.5,1.0,1.0$. During fine-tuning, we use both $\ell_{\text{dist}}$ and $\ell_{\text{motion}}$ with weights of $0.5$ and $1.0$, respectively. For each motion sequence, we crop the 3D scene that is 2.0 meters away from the root joint of the last observed pose, and the root joint is used as the origin of the cropped scene. To compute the per-vertex signed distance, we follow [38] to use 67 vertices on the human body. For the basis point set, we select 150 points that are 2.0 meters away from the origin of the scene, and those points are fixed for different sequences. More implementation details can be found in the supplementary material. We will release our code upon acceptance.
@@ 5 Conclusion
In this paper, we have introduced a framework for scene-aware human motion forecasting that explicitly models the human-scene interaction. To constrain the whole-body motion, we have proposed a mutual distance representation: the per-vertex signed distance and the per-basis point distance. The mutual distance captures both the detailed local body movements and the global human motion. Thanks to the explicit mutual distance constraints, our method yields more plausible and accurate future human motions than the state-of-the-art scene-aware motion prediction methods.
Limitations & Societal Impacts. One limitation of our work is that the motion prediction results depend on the prediction of the mutual distance. As evidenced by our ablation study with the ground truth mutual distance, improving the prediction of mutual distance leads to a better motion prediction. This therefore can be one of our future directions. In real applications, a potential risk of our model is that without explicit real world physics modeling, it is possible for our model to predict future motions that have imbalanced forces. This may lead to unstable human motion when deploy to real life applications like human-robot interaction such as collision.

### v4

@@ 4.2 Metrics, Baselines & Implementation
Baselines. We compare our method with three existing scene-aware human motion prediction models: ContAware [20], GIMO [40] and STAG [27]. ContAware [20] and STAG [27] are state-of-the-art methods which first predict the contact map and then the future motion conditioning on the predicted contact map. We train their model using the official implementation on our datasets, following their original settings to predict only human joints instead of SMPL-X poses. GIMO [40] is another state-of-the-art scene-aware human motion prediction method which mainly focuses on predicting long-term future poses given the past gaze information, and the frame rate of their motion is 2Hz. The original model of GIMO predicts future 10 frames (5 seconds) given the past 6 frames (3 seconds). We slightly adapt their official code to our settings for evaluation.
Implementation details. Our models are developed using Pytorch [24] and optimized with the ADAM [13] optimizer. Initially, we train the mutual distance prediction model and the motion forecasting model separately for 40 epochs each. Subsequently, we fine-tune both models together for an additional epoch. The original learning rate is set to 0.0005 and decays by half at the 30-th epoch for both models. We train our model on a single RTX4090. For all datasets, the loss weights of the motion forecasting model ($\lambda_{1}$,$\lambda_{2}$,$\lambda_{3}$,$\lambda_{4}$) are $1.0,0.5,1.0,1.0$. For each motion sequence, we crop the 3D scene to a region that is within 2.0 meters of the root joint of the last observed pose, and the root joint is used as the origin of the cropped scene. To compute the per-vertex signed distance, we follow [38] to use 67 vertices on the human body. For the basis point set, we select 150 points that are 2.0 meters away from the origin of the scene, and those points are fixed for different sequences.
@@ 5 Conclusion
In this work, we present an approach for scene-aware human motion forecasting that explicitly models the human-scene interaction. To constrain the whole-body motion, we have proposed a mutual distance representation: the per-vertex signed distance and the per-basis point distance. The mutual distance captures both the detailed local body movements and the global human motion. Thanks to the explicit mutual distance constraints, Our method generates future human motions that are more plausible and accurate than those produced by state-of-the-art scene-aware motion prediction techniques.
Societal Impacts & Limitations. The motion prediction results depend on the prediction of the mutual distance is one limitation of our work. Our ablation study using the ground truth mutual distance indicates that enhancing the prediction of mutual distance results in improved motion prediction. This can therefore serve as a potential direction for future research. In real applications, a potential risk of our model is that without explicit real world physics modeling, it is possible for our model to predict future motions that have imbalanced forces. This may lead to unstable human motion when deploy to real life applications like human-robot interaction such as collision.
Acknowledgements This research was supported in part by the Australia Research Council DECRA Fellowship (DE180100628) and ARC Discovery Grant (DP200102274).